Displaying reports 441-460 of 1084.Go to page Start 19 20 21 22 23 24 25 26 27 End
Reports until 11:50, Thursday 19 September 2024
Lab Infrastructure General (GQuEST)
Daniel Grass - posted 11:50, Thursday 19 September 2024 (11868)
GQuEST Optics Tables Moved into Place

[Jeff, Sander, Alex, Daniel]

We raised the screws holding the breadboard connecting the two optics tables and raised the optics tables up onto thier wheels. We then moved the tables 1' North so that the pump out station would not conflict with the wall. We moved the tables as West as possible so that the vacuum beam tubes won't hit the door frame but not so far West that the mobile clean rooms can't fit around them in their final configuration. We ensured the optics table was parallel with the West wall.

We then lowered the tables. The joint between the two tables is not level, so we will need to shim the legs (or replace the bottom of the legs). Ideally, the tables are in the same plane and all the legs make contact with the tables to prevent any rocking modes. Rigidly connecting the tables seems to remove most of these rocking modes. The gap between the table and legs is smaller than in the previous configuration.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:57, Tuesday 17 September 2024 (11867)
Moving Vacuum Chambers into B111B Progress and Future Plan

[Jeff, Ian, Torrey, Sander, JC, Lee, Daniel]

We moved a central vessel and 2 bend cubes from B150 into B111B. The central vessel is on a borrowed furniture dolly and the bend cubes are on a lowered "side table" (not an optical table).

We also lowered and moved a 2nd side table into the mobile clean room to hold and "operate on" the end cubes that are designated for the LFC.

Lastly, we raised the screws on the sled connecting the optical tables and engaged the caster wheels.

 

Jeff and I propose the future steps are as follows with a rough timeline for the rest of the week

Move the optical tables into place 1' North so that the transmitted arm pumpout stations don't hit the wall (Wednesday at 10 AM)

Maybe we move the mobile clean rooms 6" North

Place the central vessel into place (Thursday morning)

Move LFC Cubes onto side table (Thursday afternoon)

Clean B111B, with special attention to the mobile clean rooms

Open the bend cube insides to evaluate whether we need to replace the bottom flange with another flange that can hold optical posts

Replace the end cube bottom flanges with flanges with a 1/4-20 hole pattern and add my custom base plate

Place the end cubes into their LFC final position (or ~0.5" further apart than their final position and we push them together when making the vacuum system)

(Maybe replace the bend cube bottom flanges with flanges with a 1/4-20 hole pattern and) add my custom base plate

Place the bend cubes into position

Move shelf into place

Move equipment supporting optics in B102 like Mokus

Move optics sleds

Lab Infrastructure General
Daniel Grass - posted 17:16, Monday 16 September 2024 (11866)
Assembled PTA281 Optical Table Shelf in B111B

[Jeff, Sander, Torrey, Daniel]

We assembled the PTA281 Optical Table Shelf in B111B. We plan on putting this East by the end of the optics table. We decided on the shelf heights so that the IFO arms can fit under them. The table is not in its final location in the photo.

Images attached to this report
Vacuum Systems General (GQuEST)
Jeffrey Wack - posted 14:41, Monday 16 September 2024 (11865)
Cleaning of Vacuum Equipment in B150

[Daniel, Jeff]
 

We performed a rough cleaning (dusted with compressed air, wiped with isopropanol wetted Kimwipes) of the two end cubes in B150 as well as one of the service vessels. They are ready to be moved into to main lab space (B111B).

 

Images attached to this report
Computing General
Jeffrey Wack - posted 16:13, Saturday 14 September 2024 (11864)
labutils repo hosted on GitHub

I have created a labutils repo on github to be used when the McCuller Gitlab is down.

GQuEST Output Filter Cavities
Torrey Cullen - posted 16:29, Thursday 12 September 2024 (11863)
Cavity Earmuff V1

The filter cavities have very bad audio pick up, most likely coupled through sound moving the flexture mounts/mirrors. In an attempt to mitigate this we've floated the idea of cavity earmuffs. This is the first iteration of it. The design fits well enough. A few issues on this first iteration:

-We should have it contacting the metallic section of the cavity to form a seal.

-A major oversight, this was designed with only the location pictured in the above photo. This is uncharacteristic to the other 3 locations where we would need one in terms of space. For example, its never fitting here without a major redesign, or spacing out some optics.

-It would be good to glue some metal to the outside and some kind of damping plastic/foam to the inside. Will worry about this when I have a more concrete design.

I observed a spectrum of the error signal while the cavity was locked with and without the earmuff while making a similar amount of noise, and the earmuff had little to no effect. This is expected still as I only had it on one side of the cavity. Attached is the stl for this design (the print ready .3mf file for the new bambu labs printers won't upload, if you want it just reach out). Will update when I come up with a new version.

Images attached to this report
Non-image files attached to this report
Computing General
Lee McCuller - posted 10:00, Thursday 12 September 2024 (11862)
Added a lab DNS resolver for our equipment.

The NAS is running a docker instance of pi-hole That software blocks ads, but it also allows one to set up local DNS. The router should now be exposing the NAS/pihole DNS and itself as a DNS server. You can check if it is working by name resolving mokupro1 or checking that your DNS is set to 192.168.50.94 (the NAS) and 192.168.50.1 (the router).

See how to log in to add DNS entries and the password at:

https://wiki.mccullerlab.com/Main/LabSecrets

The VPN isn't yet set up to forward the DNS entries, but I'll do that soon and post a comment.

GQuEST Output Filter Cavities
Torrey Cullen - posted 15:39, Tuesday 10 September 2024 (11861)
Filter Cavity Lock Via AOM

[Jeff, Torrey]

We had some plans for filter cavity diagnostics to be done, first step in accomplishing this is locking a filter cavity with an AOM. This has been done. This post will describe the process in case anyone wants to replicate this in the future.

Attached is a screenshot showing the multi instrument mode used to accomplish this. Inputs and outputs are as follows:

Input 1 - Cavity REFL PD

Input 2 - Cavity TRANS PD

Out 1 - To the DC modulation port of the laser

Out 2 - RF in of AOM

Out 3 - RF to EOM to create sidebands

The idea is very similar as in previous posts. Create sidebands on EOM, demod in laser lock box to give an error signal, fast controller output goes to Waveform Generator which acts as a VCO (accomplished by choosing frequency modulation and input A as the reference voltage). And thats it! Relatively easy once we had the proper loop in mind. The output on the slow controller is to the laser (UGF ~10 Hz) to control any low frequency laser drifting which is known to occur. One other note that is worth mentioning is the value for the frequency modulation depth needs to be manually set. Since the FSR of the cavity is 125 MHz we initially set this range to be +/- 50 MHz to have enough range. You could also set this lower, to say 5 MHz, and then scan around on the laser to find the resonance in this range.

Images attached to this report
Equipment Loans (GQuEST)
Daniel Grass - posted 11:25, Monday 09 September 2024 (11860)
OFC Test Chamber from LIGO in B111A

[Jeff, Torrey, Daniel]

LIGO (Maty Lesovsky was my point of contact) has loaned a 24" long, 8" diameter (with 10" CF Flange) cylindrical vacuum chamber (LIGO DCC D2300310) for use in testing the output filter cavities in vacuum. Jeff and I placed it in B111A by the South optics table. I have a design for placing the OFC in the vacuum chamber. I believe I need to get a custom parts (assembly is LIGO DCC D2300308) that LIGO has designed but not procured. I will share my designs no later than next week.

Images attached to this report
Electronics General
Jeffrey Wack - posted 18:02, Thursday 05 September 2024 (11859)
Some laser modulation noise tests
coupling impedance present in pdh error?
AC 50Ohm -83dBm
DC 50Ohm -68dBm
AC 1MOhm -86dBm
DC 1MOhm -86dBm
Unplugged not present

I connected the laser DC modulation port to the Moku input of a spectrum analyzer. The spectrum analyzer always sees the 875Hz if it is connected to the laser, but the amount that the 875Hz shows up in the pdh error spectrum changes a bit on changing the spectrum analyzer input properties. Could the problem have something to do with impedance matching? the input impedance to the laser modulation port is 1kOhm.

We should plug the laser modulation port into a non-Moku spectrum analyzer and see if the problem persists.

GQuEST General (GQuEST)
Daniel Grass - posted 20:39, Wednesday 04 September 2024 (11858)
Optical Tables in B111B Joined and Moved to Final Location

[Sander, Jeff, Torrey, Daniel]

We moved the optics tables together and joined them with a 2' x 4' breadboard. The tables are unfortunately not level, likely due to the floor not being level. We think this can mostly be fixed by placing shim stock under some of the table legs.

The tables should be in their final location.

See attached photo.

Images attached to this report
Controls General
Jeffrey Wack - posted 17:23, Tuesday 03 September 2024 (11856)
775 Piezo Lock on Cavity #2

[Jeff, Torrey]

Connected the large ThorLabs piezo mounted on cavity #2 directly to the Moku (no HV amplifier).

Transmission voltage -10.7mV (no light) -9.5mV (on resonance). The photodetector in transmission was selected to measure ringdowns, it is not rated for light below 800nm. Responsivity in amps/Watt is less than 0.1

Max error amplitute is 4mV pkpk, and locking with the piezo results in constant flashing, saturating the error signal.

What voltages are we giving the piezo? The Moku has output limits of 10V. Since the maximum amplitude of the error signal is 2mV, the maximum gain we can have in our controller before saturating the Moku output is 20log_10(10/0.002) = 74dB. This is definitely being exceeded at low frequencies, we will try the HV amplifier.

The piezo for cavity control is now routed through the HV amplifier. We are able to achieve better locks but still no where sufficient for stable cavity operation. The error signal is being saturated in both directions, same as before. After some trouble shooting, two things are definitively contributing to this poor lock. The fans being on causes a non-trivial amount of noise for the piezo to handle. Additionally we found that the 840 Hz known to be coming from the poor grounding on the DC modulation port of the laser is interfering somehow. With the HV source in play we can scan and lock fully on the piezo. Getting rid of the fans and this noise introduced by having a cable plugged in into the laser seems to give a quality of lock similar to that of locking only with the laser.

 

 

Images attached to this report
GQuEST SHG (GQuEST)
Daniel Grass - posted 16:59, Tuesday 03 September 2024 (11857)
AOM Path 3 Twice Shifted Beam That Doesn't Move

[Torrey, Jeff, Daniel]

Continuing work here, we maximized the power of the beam that's shifted twice by the AOM by changing the tip and tilt of the curved mirror. We are fairly confident that this is the correct beam because it doesn't move (very much) on the camera (~6" downstream of the AOM) when we change the input frequency of the AOM by 30 MHz. The power of this beam is 0.87 mW. I then aligned this beam into a fiber collimator and got a power of 0.73 mW. We then swept the AOM frequency to evaluate how much power still gets into the beam.

Below is a table with some data. The Assumed Output Power uses the AOM specs here. It assumes for reference a 0.87 mW power for 200 MHz and is proportional to the square of the efficiency.

AOM Input Frequency (MHz) Assumed Twice Shifted Output Power (mW) Measured Twice Shifted Power Through Fiber (mW) Fiber Efficiency
200 0.87 0.73 84%
170 0.60 0.6 100%
185 0.82 0.69 84%
215 068 0.61 90%
230 0.60 0.48 80%

 

I definitely don't trust the efficiency above ~85%, but I think this gives a very good indication that we can move the AOM frequency by ~60 MHz and still have enough light to lock the cavity. 0.48 mW would saturate the fast PD, so we need to use an ND filter anyway.

I think we should copy this scheme for future AOM setups and maybe go back and adjust the mirrors. I should note that I didn't do any fine adjustment of the location of the curved mirror whatsoever. Maybe we got lucky, but this hopefully means that the path is fairly easy to set up.

The next step is to verify that this is twice shifted light. I think the easiest way is to use it in a cavity.

GQuEST SHG
Daniel Grass - posted 17:24, Friday 30 August 2024 (11855)
AOM Path 3 Initial Alignment

[Torrey, Daniel]

We aligned the beam through a f = 150 mm lens, into the AA AOM, through an iris, through a quarter waveplate, and reflected back using a concave mirror (R = 100 mm). This use of the curved mirror intended to allow us to adjust the AOM frequency without needing to realign the beam into the fiber. We currently see 3 beams from the AOM after the PBS. We aren't sure which beam is which. In theory, the twice shifted beam will not move when we change the AOM frequency. In practice, this might not be the case. We will work on this next week.

Lab Infrastructure General
Ian MacMillan - posted 14:29, Thursday 29 August 2024 (11853)
Uploading Video for Zoom Room

Upoading files for display in confrence room

Non-image files attached to this report
Atoms General
Briana Chen - posted 16:20, Wednesday 28 August 2024 - last comment - 17:17, Sunday 06 October 2024(11840)
Pump updates as of 8/22

[Briana, Ian, Torrey]

Ian suggested that one reason why the tank circuit didn't work was because the capacitance is actually 12 not 14 picofarads based on the spec sheet that came in the EOM box (linked in a previous log post). This changes the inductance from 139 nH to 119 nH, which may be within the tolerance of the inductor we used but might still amplify the wrong frequency.

Thanks to Ian and Torrey for realigning the pump before I destroyed it. We are trying to mode-match the pump and probe beams through the vapor cell (concentrate more pump power to sharpen the dip), so using JamMt and beam profiling, two lenses (150 mm and 100 mm) were added to the pump path to reduce the beam size down from ~1600 microns to ~1300 microns. One issue is that once the light passed through the EOM, the beam profile began to look more clipped, so the x and y beam diameters were not the same and changed drastically. I think this is because the beam passing through the EOM is not constant in the z-direction, so more fine-tuning is required that I was not able to finish in time. By tuning the position of the EOM, the beam size can change significantly, so it is almost certainly the EOM alignment causing this issue. Currently, from the beginning to end of the vapor cell, the probe beam has a (x,y) beam profile of (1349, 1455) and (1405, 1545) microns respectively. For the pump, this was (1368, 1361) and (1383, 1386) microns at the beginning and end of the vapor cell respectively. 

Spent some time varying the slow controller and there exists a peak around 40 Hz introduced by the locked laser (potentially due to the controller but unsure otherwise): example_peak_pt2.pngstrangepeak.png show some different controller parameters. If you reduce the unity gain frequency and the gain, this peak gets damped down, which makes sense because the gain will be lower at that frequency. This peak could be a result of us driving the system at this frequency that is in resonance with something in the system, kind of like when you have a spring and you drive it at the right frequency to increase the amplitude but once you are at higher/lower frequencies to this peak you have a lower amplitude.

We achieved fast actuation by changing the current controller. We previously did not implement this because we thought power fluctuations might become too significant. We also failed to do so initially because we needed to amplify the signal to the current controller by +14 dB in the Moku settings. On 8/21, we took data with the fast controller but the error signals looked extremely bad: horrendous_dip.pngplateau_errorsignal.png. Although the calibrated error signals for the 8/22 data also looks ugly, this one is worse because the error signal strangely plateaus. After plugging in a terminator to the BNC cable and unplugging, this plateau no longer occurred. This points to the potential fluctuations in the current controller due to the new connection from the MOD OUT port with the BNC cable. Still, the data from 8/21 shows that the slow controller is doing something. By locking and then unlocking the slow controller, we get the noise spectrum on the right side of this image: slowvsfast.png. The following are the settings used to take this data: settings.png. The controllers used were shown here: slowcontroller.pngfastcontroller.png.

The following is the error signal used for the 8/22 data. When the signal is not calibrated, it looks pretty smooth: uncalibratederrorsignal.png. Upon calibrating, the overall error signal looks bad: uglyerrorsignals.png. This is not because of the sample acquisition rate. I am pretty sure this is because the current value may be fluctuating now that we have it connected to the BNC cable/Moku. What we can do to test this is to plug in the current output to a Moku input, convert that voltage reading to current, and perform the interpolation for current similar to the one we do with temperature using the laser specs. This calibrates using a potentially varying current value along with the varying temperature value. Doing this would require three data loggers though, which is not super efficient.

On 8/22, we retook data after fixing the error signal plateau-ing. We obtain the following noise spectra: Probe(withoutpump).pdfProbe(withpump).pdf, which matches what we expect with a lower noise in the locked laser. The controllers used to lock are: slowcontroller_822.pngfastcontroller_822.png. After calibrating with the error signal (which gave us slopes of ~3.4 V/THz and ~21.6 V/THz for the probe and probe with pump respectively, showing the improved frequency discrimination), we get the following noise spectra: NoiseSpectra.pdf. This seems good but we should keep the following in mind.

Some other to-dos:

Images attached to this report
Non-image files attached to this report
Comments related to this report
Briana Chen - 17:17, Sunday 06 October 2024 (11883)

Fixed issue with terrible calibration: the temperature scan signal was weirdly noisy (see CorrectedTempscan.jpg), so I did a linear fit to smooth it over and used these values to perform calibration. This led to updated calibration values for the slope of the linear region: Probe: Slope (V/nm):  -1.659, Slope (V/THz):  3.415 | Probe with pump: Slope (V/nm):  -13.004, Slope (V/THz):  26.2367

Attached are the fixed calibration graphs: err_signal_with_pump.pdferr_signal_without_pump.pdf. The updated noise spectra with these slopes is found here: NoiseSpectra.pdf

Images attached to this comment
Non-image files attached to this comment
GQuEST Output Filter Cavities
Torrey Cullen - posted 11:09, Wednesday 28 August 2024 - last comment - 10:02, Thursday 29 August 2024(11847)
OFC1 first contact update

Cavity alignment was not recoverable. I most likely spent way too much time trying to recover instead of just starting from scratch. This is frustrating for future first contact cleaning if this cleaning procedure does improve the cavity finesse. The proccess I tried was the following:

1. Assume input alignment was unchanged, place both input couplers back on the cavity. Adjust input couplers to align both light paths back on their refl PDs. With how small the diode is on the high BW REFL PDs this should ensure 2 mirrors in their original position.

2. Piezo mirror (775 output coupler) is static and cannot be changed. Nothing to be done there.

3. Scan around on third adjustable mirror and look for any second pass beams on the camera. 

It seems to me like this should work, but this did not yield anything.I believe the piezo mirror is at a drastically different angle, which is strange because when I was doing piezo science before, changing out the piezos did not drastically change the intercavity alignment. I started over by changing the input alignment of the 1550 light and adjusting the cavity mirrors to match the 3D printed alignment tools. Doing this instead I found cavity flashes within an hour. Realigned the REFL and TRANS light onto the cameras/PDs and fine tuned alignment. OFC1 1550 light is now aligned. With the newly built tank circuit the error signal on the 1550 path looks very clean. I believe the frequency of the modulations is not quite correct and can be looked into further. I have taken ringdown measurements to see if cleaning was successful, but am getting finesse numbers greater than the specs of the mirrors. I think this may be due to the TRANS pd being saturated? Going to retry. Also need to double check the 1550 TRANS PD is fast enough, although for the 1550 light the decay time should be on the order of 3300/pi/c*2.4 = 8e-6 s or 100's of kHz. I think all the PDs we have should be this fast.

Additionally, i took off the 775 input coupler and tried to match the input alignment. As built, the 775 light first sees the piezo mirror after the input coupler. For any given input alignment the piezo mirror is putting the beam into the metal casing of the cavity. Need to explore this still.

Comments related to this report
Lee McCuller - 10:02, Thursday 29 August 2024 (11849)

Wow, that's awesome to hear. Good optics and good procedures.

Torrey Cullen - 14:18, Wednesday 28 August 2024 (11848)

Update:
Retook the measurement, making sure to move the laser frequency away from equilibrium enough via the output offset, so that when I break lock no new light is resonant and I am purely measuring the light exiting the cavity. Additionally reduced input power slightly so that the PD is not saturated. Did two measurements in transmission of the cavity and find finesse measurements very close to the specs of the mirror. I think more analysis is needed (align 775 and do similar measurements) but we may be able to conclude that the mirrors were simply dirty coming from the manufacturer. Plot attached. It seems like this may be worth doing for the second cavity.

Images attached to this comment
GQuEST Output Filter Cavities
Torrey Cullen - posted 14:47, Wednesday 21 August 2024 - last comment - 13:09, Thursday 29 August 2024(11845)
OFC1 Super optic cleaning

Super optic cleaning prep notes:

I've moved the bin of extra OFC parts to the other table in order to clear room on the OFC table. This will allow us to keep the exposed super optics inside the clean room at all times. I've set up a prep area on the OFC table closest to the door to B102. I have cleaned a pair of metal tweezers (the plastic ones in trials were very staticy), all of the required hex keys, and moved the first contact supplies to this location.

Currently my plan is to fully remove the ~2 inch spacer/flexture mount combination along with the optic. It would be possible to unscrew the nylon tip set screw that holds the mirrors in place but this seems like it adds unneeded risk and the flexture mount combinations can be removed/replaced and still retain inter-cavity alignment.

The piezo mirror must be fully removed in order to access the mirror. I have set aside an SM1 ring spanner wrench and piezo alignment tool to be cleaned when I am garbed up.

Two of the mirrors on one side (OFC1:FCM2 and OFC1:FCM4) are too close to a beamsplitter (OFC1:BS2) to remove. This BS must be removed from the table temporarily. I have done this.

End of preparation for first contacting particle count 0/0/0 and 1/0/0 during two 60 second count trials.

** Lunch**

Midway particle count: 0/0/0

Plan is to clean both the HR and AR sides. I think the way to do it is rest the optic on the AR side, paint the HR side, let this side dry, flip it and paint the AR side while the optic is resting on the painted HR side, wait til both are dry, remove both sides without resting the optic and install back into mount.

HR side first contact complete. Piezo mirror was visibly dirty. Looked like metal shavings. The AR side also had a large imprint on it from the piezo being driven a couple of times. Similar to what we've seen in the past.

AR side first contact complete. Mirrors are resting on dry HR side. Waiting 10 more minutes and then peeling.

Cleaned the inside circular part of the spacers that hold the mirror. The piezo mirror mount one had what looked like small metalic particulates.

All 8 peels were very successful. I noticed some specs (most likely stringing from the FC) on the AR side of OFC1:FCM1. Repainted this side while it was in its mount. Peel was successful. Reinstalled the mirrors on the cavity and now will scan and look to realign.

End of work particle count: 0/0/0

 

Comments related to this report
Torrey Cullen - 13:09, Thursday 29 August 2024 (11850)

Forgot to provide photos taken during the cleaning. See attached.

Images attached to this comment
Torrey Cullen - 15:15, Thursday 22 August 2024 (11846)

It seems like the alignment is not recoverable. Tried a few tricks to quickly regain alignment but it has not worked. Will need to start from scratch. 

Lab Infrastructure General (Layouts)
Sander Vermeulen - posted 19:44, Tuesday 20 August 2024 - last comment - 13:15, Thursday 29 August 2024(11844)
Moved optical tables in B111B

[Jeff, Daniel, Sander]

We have moved the TMC optical tables into close to their 'final' positions, according to the designed lab layout (see attached). We extended the caster wheels in the tables, which allows them to be moved around with two people pushing and one person guiding the table.  The east table is not pushed up to the other table fully. We need to find some way to do this without bumping the tables into eachother and end up with both tables in the final desired positions. Note that the table tops are raised when the wheels are extended (vs. wheels retracted) and that the breadbord table tops protrude slightly from the base of the top (i.e. there's a horizontal 'lip' at the top), which makes this more difficult. 

Some observations:

- As mentioned previously, the tabletops are not rigidly attached to the legs. This means the tops can and will move slightly on the tops of the pistons (which are on top of the legs) that the tabletop rests on, limited only by friction. The only thing preventing the top from moving completely off the legs are the earthquake stop bars.

- It seems the heads of the pistons (black part in photo), are not aligned with the bodies of the pistons (silver part in photo); inadvertently moving the top w.r.t the legs probably caused this misalignment. This misalignment is quite bad for some of the legs (see photo). 

- The table tops do not rest equally on all four legs, there are visible gaps between the tops of the pistons and the top (see photo). This is probably because the lab floor is not level. 

I think we can do the final movement by extending wheels on both tables, so that the breadbord lips are at the same height and the tables can recoil when pushing them together. Then we can try to slowly push the tables together so there's no gap, which will likely move the tables away from their desired final position. Then we could put e.g. a breadbord over the seam to make a rigid connection, after which we can do the final positioning, moving both tables at the same time. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 13:15, Thursday 29 August 2024 (11851)

You have the SHG sled and power distribution on this diagram twice just fyi.

GQuEST Output Filter Cavities
Daniel Grass - posted 00:34, Friday 16 August 2024 - last comment - 09:48, Friday 30 August 2024(11838)
Differential Length Fluctuation of 775 nm light vs. 1550 nm light due to ThermoOptic Effects in the GQuEST Output Filter Cavities

We are currently wondering why the 1550 nm light is not co-resonant with the 775 nm light, used to lock the GQuEST Output Filter Cavities, over time. Some ideas are the AOM frequency drifts (unlikely), but a potential culperate is differential length changes of the cavity due to thermooptic effects (thermal expansion and thermo-refraction). For light that hits the surface of the coating, Evans (2008) (Eq. 3) gives the sensitivity of the sensed position of the mirror \Delta T to temperature fluctuations T as

\[ \frac{\partial \Delta z}{\partial T} = \bar{\alpha}_c h_c - \bar{\beta}\lambda \]

Where \bar{\alpha}_c is the effective coefficient of thermal expansion, h_c is the coating thickness, \bar{\beta} is the effective thermo-refractive coefficient, and \lambda is the wavelength. This expresion doesn't work for the 775 nm light since it is below a layer of 1550 nm coating. Assume a fraction (\gamma) of the coating is for 1550 nm light and (1-\gamma) of the coating is for 775 nm light. \gamma \approx 0.5, although the 775 nm light coating is probably thinner since the fractional wavelength stacks are thinner. This analysis might also not be exact for the mirrors since none of them are HR for both wavelengths.

\[ \frac{\partial \Delta z_{775~\text{nm}}}{\partial T} = (1-\gamma)\bar{\alpha}_c h_c - \gamma\bar{\alpha}_c h_c(n-1) - \bar{\beta}( \gamma h_c+\lambda) \]

Where n is the index of refraction of 775 nm light in the 1550 nm coating. There is a minus sign in front of it because a higher index causes more phase to be accumulated than in air.

Now consider the differential sensitivity of the sensed position of the mirror, assuming the thermorefractive coefficient is very roughly the same for both wavelengths.

\[ \Delta \frac{\partial \Delta z}{\partial T} \equiv \frac{\partial \Delta z_{1550~\text{nm}}}{\partial T} -  \frac{\partial \Delta z_{775~\text{nm}}}{\partial T}\]

\[ \Delta \frac{\partial \Delta z}{\partial T} = \gamma h_c (\bar{\alpha}_c n + \bar{\beta}) \]

Plugging in numbers from our GQuEST Paper Table II, 

\[ \Delta \frac{\partial \Delta z}{\partial T} \approx 2 \cdot 10^{-10}~\text{m/K} \]

I'm not sure if ~1/1000 of a wavelength can explain the drift in the co-resonance over time.

Comments related to this report
Daniel Grass - 09:48, Friday 30 August 2024 (11854)

Using Delta f /f = Delta l / L, where f is the frequency of the light and L is the length of the cavity, we get 16 kHz / K, not enough to explain the overnight drift of ~ 1 MHz. Lee points out that this temperature dependence might be enhanced by the quarter and half wavelength stacks.

Displaying reports 441-460 of 1084.Go to page Start 19 20 21 22 23 24 25 26 27 End